MILLIMETER/SUBMILLIMETER SPECTRUM OF THE CCP (X^{2}Π_{r}) RADICAL

\underline{D. T. HALFEN}, Steward Observatory, University of Arizona, Tucson, AZ, 85721; D. J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY, 40506; and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721.

The pure rotational spectrum of the CCP radical (X^{2}Π_{r}) has been measured using millimeter-wave direct absorption techniques. This work is the first high-resolution data for this radical. This species was created by the reaction of gasphase phosphorus and acetylene in the presence of argon carrier gas and an AC glow discharge. Twenty-four rotational transitions were measured in the frequency range of 120 to 413 GHz. Each transition is separated into two spin-orbit components; however, only the Ω = 1/2 component has been observed to date. This spin component is additionally split into lambda-doublets. Below 209 GHz, hyperfine interactions due to the phosphorus nuclear spin of I = 1/2 are evident, spitting each lambda-doublet into two lines. The data were fit with a case (c) Hamiltonian, and effective rotational, lambda-doubling, and phosphorus hyperfine constants were determined. Additional searches are being conducted for the Ω = 3/2 component. These data prove that CCP has a linear geometry in its ground state. This structure is similar to other MC_{2} species, such as CCS and CCCl, in contrast to NaC_{2}, MgC_{2}, AlC_{2}, and SiC_{2}. The change in structure from T-shaped to linear is caused by a increase in the covalent bonding interactions between the C_{2} moiety and the heteroatom.